OpenStudy (anonymous):

What is true about the solutions of a quadratic equation when the radicand of the quadratic formula is a perfect square? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (campbell_st):

rational, and they can be unequal, a trinomial or equal for a perfect square

OpenStudy (anonymous):

so its C?

OpenStudy (anonymous):

say \(b^2-4ac=m^2\) \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}=\frac{-b\pm\sqrt{m^2}}{2a}=\frac{-b \pm m}{2a}\] two rational numbers unless of course \(m=0\) which is also a perfect square but seems to be ignored in this question

OpenStudy (campbell_st):

thats would be my choice

OpenStudy (anonymous):

yeah mine too, even though strictly speaking it is wrong

OpenStudy (anonymous):

lol alright thank you guys

OpenStudy (campbell_st):

I agree... as a perfect square will have a square number for the discriminant....

OpenStudy (anonymous):

so what do you mean.. its B?

OpenStudy (campbell_st):

no its C but a poorly worded question